If cost varies inversely as the square of the number purchased, and 5 can be purchased for $45, how much would 10 cost?
Answer by mukhopadhyay(490)

Cost = k/(number)^2
So, 45 = k/25
=> k = 45*25
Cost for 10 would be
Cost = 45*25/10^2
=> Cost = 45*25/100 = 45/4 = $11.25